# GMT: Guided Mask Transformer for Leaf Instance Segmentation

### v2

@@ Abstract
Leaf instance segmentation is a challenging multi-instance segmentation task, aiming to separate and delineate each leaf in an image of a plant. Accurate segmentation of each leaf is crucial for plant-related applications such as the fine-grained monitoring of plant growth and crop yield estimation. This task is challenging because of the high similarity (in shape and colour), great size variation, and heavy occlusions among leaf instances. Furthermore, the typically small size of annotated leaf datasets makes it more difficult to learn the distinctive features needed for precise segmentation. We hypothesise that the key to overcoming the these challenges lies in the specific spatial patterns of leaf distribution. In this paper, we propose the Guided Mask Transformer (GMT), which leverages and integrates leaf spatial distribution priors into a Transformer-based segmentor. These spatial priors are embedded in a set of guide functions that map leaves at different positions into a more separable embedding space. Our GMT consistently outperforms the state-of-the-art on three public plant datasets.
@@ 1 Introduction
Instance segmentation is a well-studied computer vision task aiming at precisely delineating each object within an image at the pixel level [15, 32]. While state-of-the-art (SotA) Transformer models [40, 12, 4] have outperformed convolutional neural networks (CNNs) in various tasks, they still struggle to achieve robust instance segmentation for plant leaves. This challenge is caused by the small size of most plant datasets, which typically contain only a few hundred to a few thousand images, making it difficult to train large models effectively. Moreover, plants exhibit severe overlaps and a huge intra- and inter-species morphological variability [36], which further challenges computer vision approaches. As illustrated in Figure 1, current Transformer-based segmentors such as Mask2Former [8] (column 3) often fail to segment leaves correctly, missing small and overlapping leaves (row 1) or confusing other green objects with leaves (row 2). The challenges arising from leaf instance segmentation have been extensively studied for years in the CVPPP/CVPPA series of workshops. [^1] https://cvppa2021.github.io/challenges/
@@
In many plants, particularly those with a rosette leaf growth pattern (e.g. Arabidopsis), leaves grow outward from the center, with those near the center being younger (typically smaller and more densely clustered), while those farther out are older and tend to be larger [14]. Also, leaves at different growth stages exhibit distinct shapes. In other words, leaf positions are correlated to the challenges in plant leaf instance segmentation. These observations motivate us to incorporate leaf position as prior knowledge to improve segmentation models. Therefore, we present the Guided Mask Transformer (GMT), a Transformer-based model tailored to instance segmentation on plant images (Figure 2). Our model introduces three novel components, namely Guided Positional Encoding (GPE), Guided Dynamic Positional Queries (GDPQ), and Guided Embedding Fusion Module (GEFM), to integrate prior knowledge on the spatial distributions of leaf instances which is represented by guide functions. Specifically, GPE and GDPQ enhance the positional representation of instances within the attention mechanism, while GEFM improves the separability of feature representations for different instances.
@@
Experimental results on three popular plant datasets [33, 10, 38] demonstrate that GMT outperforms SotA leaf instance segmentation models. Quantitative and qualitative results show that the performance improvements are most notable for small to medium-sized leaves, especially for the overlapping ones (Figure 1 top right), and the confusion with leaf-like objects is reduced (Figure 1 bottom right). This highlights our method’s ability of addressing the most challenging scenarios in plant analysis. We also present ablation studies for GMT’s key components, demonstrating that its overall design is crucial to the learning process.
In summary, the key contributions of this work are:
  • We introduce GMT, a Transformer-based segmentor that effectively incorporates leaf-position-related prior knowledge to improve leaf instance segmentation.
@@
  • We show SotA performance of GMT on three public datasets and discuss its ability to tackle the most challenging scenarios in plant data analysis, with ablation studies validating the importance of its overall design.
@@ 2 Related Work
We review plant image segmentation approaches and, since our approach is based on Transformers, we also review recent works in Transformer-based segmentation.
@@ 3 Proposed Method
Our main idea is to first capture and represent the unique characteristics of leaves and plants, and then devise an effective mechanism to integrate these characteristics into a segmentation model. We utilise a set of guide functions to represent these characteristics. These guide functions are adjusted to the pixel coordinates of different instances and optimised to separate distinct instances within an embedding space. Subsequently, the proposed Guided Mask Transformer (GMT), depicted in Figure 2, leverages the optimised guide functions, which possess instance distribution priors, to train predicting instance masks of plant leaves.
@@ 3.1 Guide Functions
After training, we obtain $d_{g}$ guide functions, each with its own learned parameters $\psi_{i}$ (c.f. Equation 1). These guide functions are adjusted to the pixel coordinates of different instances (c.f. Equation 1), allowing them to carry prior information of instance locations; they are also able to embed distinct instance masks and separate them in the embedding space (c.f. Section 3.1).
@@ 3.2 GMT: Guided Mask Transformer
Once the guide functions are obtained, we incorporate them into the proposed GMT model. Our key contribution lies in the novel modules we introduce, GPE, GEFM, and GDPQ, to fully leverage these functions within a Transformer-based architecture, as detailed at Section 3.2.2. We extend the meta-architecture of Mask2Former [8] with our proposed modules, given its strong instance segmentation capabilities in general domains but limitations in leaf instance segmentation (c.f. Figure 1 column 3).
@@ 3.2.1 Mask2Former Preliminaries
where $V\in\mathbb{R}^{hw\times d_{p}}$ is identical to $K$, and the first term is known as the cross-attention map, $A=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{p}}}\right)$. Equation 5 shows that cross-attention aggregates image context based on the dot-product similarity between the queries $Q$ and keys $K$, Both the content and positional components of $Q$ and $K$ contribute to the attention scores, leveraging both semantic and spatial information.
@@ 3.2.2 Guided Mask Transformer
Our GMT follows the standard design of Mask2Former, using MSDeformAttn [48] as the pixel decoder and a Transformer decoder composed of nine blocks, each with a masked cross-attention layer, a self-attention layer, and a feed-forward layer. Intermediate supervision is applied after each Transformer block predicting instance masks. Below we describe the key components of GMT, which replace or enhance parts of the pixel decoder and Transformer decoder. These adaptations enable the effective integration of the guide functions obtained from the previous stage. The overall architecture of GMT is shown in Figure 2.
Guided Positional Encoding (GPE). As introduced in Section 3.2.1, the positional components of pixel features in the pixel decoder enhance spatial representation of instances. We propose GPE to better align this spatial representation with the prior knowledge of leaf instance distribution.
@@
As presented in Section 3.1 and Equation 1, the guide functions are sinusoidal, correspond to the pixel coordinates of instances, and possess positional priors of leaf instances after training. This makes them suitable for improving the spatial representation of the original SPE. However, the number of guide functions $d_{g}$ is much smaller than the pixel feature dimension $d_{p}$; for instance, in our experiments, $d_{g}=16$ and $d_{p}={256}$. An intuitive approach is to increase $d_{g}$ to $d_{p}$ when learning the guide functions. However, we found this to be impractical due to high computational cost and optimisation challenges for the guide functions at such high dimensionality. Therefore, we propose a “postprocessing” method to expand the low-dimensional guide functions to high-dimensional positional encodings. We expand the number of guide functions to $d_{p}$ by maintaining the frequency components $\psi{\left[1\right]}$ and $\psi{\left[2\right]}$, while shifting the phase components $\psi{\left[3\right]}$. Specifically, every guide function $f_{i}$ is expanded to $J=d_{p}/d_{g}$ functions, defined as:
@@
Figure 4: GDPQ Module. The positional queries at current Transformer block $Q^{t}_{p}$ is dynamically generated on the guide-function-encoded mask predictions from last block. $Q^{t-1}$ denotes the object queries from last Transformer block. The dimensions of different elements are shown: $n$ is the length of object queries, and $h$ and $w$ denote the spatial dimension of the final pixel features. The computation process of $\mathcal{E}(\mathcal{S}^{t-1};\Psi)$ as presented in Equation 10 is highlighted in red.
Guided Embedding Fusion Module (GEFM). As shown in Section 3.1, the trained guide functions can map the instance masks to guided embeddings where different instances are well separated. To facilitate this mapping in GMT and enhance the distinction between similar leaf instances, we use a $1\times 1$ convolution layer to project the pixel decoder outputs to the guided features (Figure 2). These guided features are supervised by the ground-truth guided embeddings (obtained by encoding the ground-truth instance masks with the trained guide functions), with an $L_{1}$ loss [^2] The loss is weighted higher on the edges of instances, following the implementation of [24]., as shown in Figure 3. Another $1\times 1$ convolution layer is used to fuse the concatenation of guided features and the original mask features (Figure 2).
Guided Dynamic Positional Queries (GDPQ). As discussed in Section 3.2.1, the Transformer decoder refines object queries $Q$, which are composed of content queries $Q_{c}$ and positional queries $Q_{p}$, through attention mechanisms to capture semantic and spatial information of instances. Typically, $Q_{p}$ are randomly initialised and updated along with $Q_{c}$. However, this is insufficient for learning the useful spatial representations of leaf instances under complex plant structures, particularly with small labelled datasets. Hence, we propose GDPQ, which dynamically generates positional queries conditioned on guide functions.
@@
where $h$ is an MLP, $A^{t-1}$ is the cross-attention maps (c.f. Equation 5), $K_{p}^{t-1}$ denotes the positional encodings of pixel features, and $B$ is a trainable bias term.
In our case, since the trained guide functions possess instance location priors, we use them to embed the intermediate predictions of GMT and generate positional queries based on the resulting embeddings, as shown in Figure 4. The proposed GDPQ is expressed as:
@@
where $\mathcal{S}^{t-1}$ denotes the set of predicted instance masks from the previous Transformer block, and $\mathcal{E}(\mathcal{S}^{t-1};\Psi)=\operatorname*{Concat}_{S\in\mathcal{S}^{t-1}}e(S;\Psi)$ represents the concatenation of all guided embeddings (c.f. Equation (2)(3)) of these masks. The computation process of $\mathcal{E}(\mathcal{S}^{t-1};\Psi)$ is also highlighted in Figure 4. We configure $h\left(\cdot\right)$ as a three-layer MLP to map the concatenated embeddings to the same dimension as object queries (from $d_{g}$ to $d_{p}$). In this manner, the positional queries are conditioned on the guide functions, which embody positional priors, and are dynamically generated based on intermediate mask predictions, which continue to refine as the Transformer decoder progresses deeper. Finally, the positional queries generated by GDPQ are summed with the content queries to form new object queries, which are then processed by the Transformer blocks.
@@
where $L_{\text{guide}}=\lambda_{\text{guide}}L_{\text{1}}$ is a weighted $L_{\text{1}}$ loss applied to the guided features, as described in GEFM and Figure 3. $L_{\text{M2F}}=\lambda_{\text{ce}}L_{\text{ce}}+\lambda_{\text{dice}}L_{\text{dice}}+\lambda_{\text{cls}}L_{\text{cls}}$ follows the standard Mask2Former loss [8], with $L_{\text{ce}}$ and $L_{\text{dice}}$ denoting binary cross-entropy loss and Dice loss [31] for mask prediction, and $L_{\text{cls}}$ as the cross-entropy loss for classification. The $\lambda$ terms denote the weights for each corresponding loss.
@@ 4.2 Implementation Details
Model Settings. By default, both the baseline model (Mask2Former) and our GMT use a ResNet-50 [18] backbone. We evaluate other backbone options (ResNet-101 and Swin Transformer [29]) on GMT at Section 4.4. As the number of training images of a leaf dataset is typically small, the baseline model and GMT [^3] For GMT, we load the COCO pre-trained weights from Mask2Former, omitting the incompatiable keys. are pre-trained on COCO instance segmentation dataset [26].
@@ 4.3 Quantitative and Qualitative Results
The overall quantitative results on CVPPP LSC, MSU-PID and KOMATSUNA are presented on Table 1, with qualitative results shown in Figure 5. Additional visual results are shown in Supplementary Materials. For all three datasets, we observe that Mask2Former (COCO pre-trained) already surpass previous SotA approaches, so we mainly compare Mask2Former with the proposed GMT.
@@
CVPPP LSC. We observe consistent performance improvements of GMT compared to Mask2Former, as shown in the first part of Table 1, with increases of +0.8 in BD, +0.6 in SBD and +0.19 in ||||DiC||||. Figure 5 row 1 demonstrates that GMT is able to delineate the shape of small leaves at the plant center more accurately.
MSU-PID. Table 1 second part shows that GMT surpass Mask2Former by +0.4 on BD and SBD, respectively, while performing on par in ||||DiC||||. Figure 5 row 2 shows that GMT mitigates under-segmentation when one tiny leaf overlaps another leaf, whereas Mask2Former merges the two leaves and misses part of the outer shape; row 3 illustrates that GMT reduces the occurrence of missing leaves.
KOMATSUNA. As shown in the third section of Table 1, GMT improves over Mask2Former with a +0.5 increase in BD and slightly outperforms it in SBD by +0.1, while showing a minor underperformance in ||||DiC|||| by -0.02. Figure 5 row 4 demonstrates that GMT avoids confusing similar green objects with leaves, and row 5, again, shows that GMT reduces under-segmentation in occlusion scenarios.
@@
The results shown in Table 3 demonstrate that for small and medium leaves on both datasets, GMT improve the segmentation quality (i.e. BD and SBD) notably while its counting ability remains on par with Mask2Former. For large leaves in MSU-PID, GMT performs worse in BD but outperforms in SBD and ||||DiC||||. As discussed in Section 4.2, BD may underestimate the errors caused by under-segmentation or missing leaves due to its one-directional matching nature, while SBD is a more balanced metric for segmentation quality. In fact, the better ||||DiC|||| score of GMT suggests that, despite a lower BD compared to Mask2Former, GMT is more effective at identifying all leaves. Finally, for large leaves in KOMATSUNA, GMT performs comparably to Mask2Former. Overall, GMT excels in segmenting small and medium leaves, which are more challenging in leaf instance segmentation, while maintaining solid performance on large leaves.

### v3

@@ Abstract
Leaf instance segmentation is a challenging multi-instance segmentation task, aiming to separate and delineate each leaf in an image of a plant. Accurate segmentation of each leaf is crucial for plant-related applications such as the fine-grained monitoring of plant growth and crop yield estimation. This task is challenging because of the high similarity (in shape and colour), great size variation, and heavy occlusions among leaf instances. Furthermore, the typically small size of annotated leaf datasets makes it more difficult to learn the distinctive features needed for precise segmentation. We hypothesise that the key to overcoming the these challenges lies in the specific spatial patterns of leaf distribution. In this paper, we propose the Guided Mask Transformer (GMT), which leverages and integrates leaf spatial distribution priors into a Transformer-based segmentor. These spatial priors are embedded in a set of guide functions that map leaves at different positions into a more separable embedding space. Our GMT consistently outperforms the state-of-the-art on three public plant datasets. Our code is available at https://github.com/vios-s/gmt-leaf-ins-seg.
@@ 1 Introduction
Instance segmentation is a well-studied computer vision task aiming at precisely delineating each object within an image at the pixel level [15, 33]. While state-of-the-art (SOTA) Transformer models [41, 12, 4] have outperformed convolutional neural networks (CNNs) in various tasks, they still struggle to achieve robust instance segmentation for plant leaves. This challenge is caused by the small size of most plant datasets, which typically contain only a few hundred to a few thousand images, making it difficult to train large models effectively. Moreover, plants exhibit severe overlaps and a huge intra- and inter-species morphological variability [37], which further challenges computer vision approaches. As illustrated in Fig. 1, current Transformer-based segmentors such as Mask2Former [8] (column 3) often fail to segment leaves correctly, missing small and overlapping leaves (row 1) or confusing other green objects with leaves (row 2). The challenges arising from leaf instance segmentation have been extensively studied for years in the CVPPP/CVPPA series of workshops. [^1] https://cvppa2021.github.io/challenges/
In many plants, particularly rosette plants (e.g. Arabidopsis), leaves grow outward from the centre, with those near the centre being younger (typically smaller and more densely clustered), while those farther out are older and tend to be larger [14]. Also, leaves at different growth stages exhibit distinct shapes. These features can be captured via top-down views of plants (examples in Fig. 1) using laboratory imaging devices or UAVs [19].
@@
In other words, leaf positions are correlated to the challenges in plant leaf instance segmentation. This correlation motivates us to incorporate leaf position as prior knowledge to improve segmentation models. Therefore, we present the Guided Mask Transformer (GMT), a Transformer-based model tailored to instance segmentation on plant images (Fig. 2). Our model introduces three novel components, namely Guided Positional Encoding (GPE), Guided Dynamic Positional Queries (GDPQ), and Guided Embedding Fusion Module (GEFM), to integrate prior knowledge on the spatial distributions of leaf instances which is represented by guide functions. Specifically, GPE and GDPQ enhance the positional representation of instances within the attention mechanism, while GEFM improves the separability of feature representations for different instances.
@@
Experimental results on three popular plant datasets [34, 10, 39] demonstrate that GMT outperforms SOTA leaf instance segmentation models. Quantitative and qualitative results show that the performance improvements are most notable for small to medium-sized leaves, especially for the overlapping ones (Fig. 1 top right), and the confusion with leaf-like objects is reduced (Fig. 1 bottom right). This highlights our method’s ability of addressing the most challenging scenarios in plant analysis. We also present ablation studies for GMT’s key components, demonstrating that its overall design is crucial to the learning process.
@@
In summary, our key contributions are: (i) we introduce GMT, a Transformer-based segmentor that leverages position-related prior knowledge to improve leaf instance segmentation; (ii) we show SOTA performance on three public datasets and discuss its ability to tackle the most challenging scenarios in plant data analysis, with ablation studies validating the importance of its overall design.
@@ 3 Proposed Method
The Guided Mask Transformer (GMT) (c.f. Fig. 2) captures and represents unique leaf and plant characteristics, integrating them into a segmentation model. Using guide functions tailored to pixel coordinates, we optimise these functions to distinguish instances within an embedding space. Then, GMT leverages these optimised guide functions with instance distribution priors to predict plant leaf instance masks.
@@ 3.1 Guide Functions
After training, we obtain $d_{g}$ guide functions, each with its own learned parameters $\psi_{i}$ (c.f. Eq. 1). These guide functions are adjusted to the pixel coordinates of different instances (c.f. Eq. 1), allowing them to carry prior information of instance locations; they are also able to embed distinct instance masks and separate them in the embedding space (c.f. Sec. 3.1).
@@ 3.2 GMT: Guided Mask Transformer
Once the guide functions are obtained, we incorporate them into the proposed GMT model. Our key contribution lies in the novel modules we introduce, GPE, GEFM, and GDPQ, to fully leverage these functions within a Transformer-based architecture, as detailed at Sec. 3.2.2. We extend the meta-architecture of Mask2Former [8] with our proposed modules, given its strong instance segmentation capabilities in general domains but limitations in leaf instance segmentation (c.f. Fig. 1 column 3).
@@ 3.2.1 Mask2Former Preliminaries
where $V\in\mathbb{R}^{hw\times d_{p}}$ is identical to $K$, and the first term is known as the cross-attention map, $A=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{p}}}\right)$. Eq. 5 shows that cross-attention aggregates image context based on the dot-product similarity between the queries $Q$ and keys $K$, Both the content and positional components of $Q$ and $K$ contribute to the attention scores, leveraging both semantic and spatial information.
@@ 3.2.2 Guided Mask Transformer
Our GMT follows the standard design of Mask2Former, using MSDeformAttn [49] as the pixel decoder and a Transformer decoder composed of nine blocks, each with a masked cross-attention layer, a self-attention layer, and a feed-forward layer. Intermediate supervision is applied after each Transformer block predicting instance masks. Below we describe the key components of GMT, which replace or enhance parts of the pixel decoder and Transformer decoder. These adaptations enable the effective integration of the guide functions obtained from the previous stage. The overall architecture of GMT is shown in Fig. 2.
Guided Positional Encoding (GPE). As introduced in Sec. 3.2.1, the positional components of pixel features in the pixel decoder enhance spatial representation of instances. We propose GPE to better align this spatial representation with the prior knowledge of leaf instance distribution.
@@
As presented in Sec. 3.1 and Eq. 1, the guide functions are sinusoidal, correspond to the pixel coordinates of instances, and possess positional priors of leaf instances after training. This makes them suitable for improving the spatial representation of the original SPE. However, the number of guide functions $d_{g}$ is much smaller than the pixel feature dimension $d_{p}$; for instance, in our experiments, $d_{g}=16$ and $d_{p}={256}$. An intuitive approach is to increase $d_{g}$ to $d_{p}$ when learning the guide functions. However, we found this to be impractical due to high computational cost and optimisation challenges for the guide functions at such high dimensionality. Therefore, we propose a “postprocessing” method to expand the low-dimensional guide functions to high-dimensional positional encodings. We expand the number of guide functions to $d_{p}$ by maintaining the frequency components $\psi{\left[1\right]}$ and $\psi{\left[2\right]}$, while shifting the phase components $\psi{\left[3\right]}$. Specifically, every guide function $f_{i}$ is expanded to $J=d_{p}/d_{g}$ functions, defined as:
@@
Figure 4: GDPQ Module. The positional queries at current Transformer block $Q^{t}_{p}$ is dynamically generated on the guide-function-encoded mask predictions from last block. $Q^{t-1}$ denotes the object queries from last Transformer block. The dimensions of different elements are shown: $n$ is the length of object queries, and $h$ and $w$ denote the spatial dimension of the final pixel features. The computation process of $\mathcal{E}(\mathcal{S}^{t-1};\Psi)$ as presented in Eq. 10 is highlighted in red.
Guided Embedding Fusion Module (GEFM). As shown in Sec. 3.1, the trained guide functions can map the instance masks to guided embeddings where different instances are well separated. To facilitate this mapping in GMT and enhance the distinction between similar leaf instances, we use a $1\times 1$ convolution layer to project the pixel decoder outputs to the guided features (Fig. 2). These guided features are supervised by the ground-truth guided embeddings (obtained by encoding the ground-truth instance masks with the trained guide functions), with an $L_{1}$ loss [^2] The loss is weighted higher on the edges of instances, following the implementation of [25]., as shown in Fig. 3. Another $1\times 1$ convolution layer is used to fuse the concatenation of guided features and the original mask features (Fig. 2).
Guided Dynamic Positional Queries (GDPQ). As discussed in Sec. 3.2.1, the Transformer decoder refines object queries $Q$, which are composed of content queries $Q_{c}$ and positional queries $Q_{p}$, through attention mechanisms to capture semantic and spatial information of instances. Typically, $Q_{p}$ are randomly initialised and updated along with $Q_{c}$. However, this is insufficient for learning the useful spatial representations of leaf instances under complex plant structures, particularly with small labelled datasets. Hence, we propose GDPQ, which dynamically generates positional queries conditioned on guide functions.
@@
where $h$ is an MLP, $A^{t-1}$ is the cross-attention maps (c.f. Eq. 5), $K_{p}^{t-1}$ denotes the positional encodings of pixel features, and $B$ is a trainable bias term.
In our case, since the trained guide functions possess instance location priors, we use them to embed the intermediate predictions of GMT and generate positional queries based on the resulting embeddings, as shown in Fig. 4. The proposed GDPQ is expressed as:
@@
where $\mathcal{S}^{t-1}$ denotes the set of predicted instance masks from the previous Transformer block, and $\mathcal{E}(\mathcal{S}^{t-1};\Psi)=\operatorname*{Concat}_{S\in\mathcal{S}^{t-1}}e(S;\Psi)$ represents the concatenation of all guided embeddings (c.f. Equation (2)(3)) of these masks. The computation process of $\mathcal{E}(\mathcal{S}^{t-1};\Psi)$ is also highlighted in Fig. 4. We configure $h\left(\cdot\right)$ as a three-layer MLP to map the concatenated embeddings to the same dimension as object queries (from $d_{g}$ to $d_{p}$). In this manner, the positional queries are conditioned on the guide functions, which embody positional priors, and are dynamically generated based on intermediate mask predictions, which continue to refine as the Transformer decoder progresses deeper. Finally, the positional queries generated by GDPQ are summed with the content queries to form new object queries, which are then processed by the Transformer blocks.
@@
where $L_{\text{guide}}=\lambda_{\text{guide}}L_{\text{1}}$ is a weighted $L_{\text{1}}$ loss applied to the guided features, as described in GEFM and Fig. 3. $L_{\text{M2F}}=\lambda_{\text{ce}}L_{\text{ce}}+\lambda_{\text{dice}}L_{\text{dice}}+\lambda_{\text{cls}}L_{\text{cls}}$ follows the standard Mask2Former loss [8], with $L_{\text{ce}}$ and $L_{\text{dice}}$ denoting binary cross-entropy loss and Dice loss [32] for mask prediction, and $L_{\text{cls}}$ as the cross-entropy loss for classification. The $\lambda$ terms denote the weights for each corresponding loss.
@@ 4.2 Implementation Details
Model Settings. By default, both the baseline model (Mask2Former) and our GMT use a ResNet-50 [18] backbone. We evaluate other backbone options (ResNet-101 and Swin Transformer [30]) on GMT at Sec. 4.4. As the number of training images of a leaf dataset is typically small, the baseline model and GMT [^3] For GMT, we load the COCO pre-trained weights from Mask2Former, omitting the incompatiable keys. are pre-trained on COCO instance segmentation dataset [27].
@@ 4.3 Quantitative and Qualitative Results
The overall quantitative results on CVPPP LSC, MSU-PID and KOMATSUNA are presented on Tab. 1, with qualitative results shown in Fig. 5. Additional visual results are shown in Supplementary Materials. For all three datasets, we observe that Mask2Former (COCO pre-trained) already surpass previous SOTA approaches, so we mainly compare Mask2Former with the proposed GMT.
@@
CVPPP LSC. We observe consistent performance improvements of GMT compared to Mask2Former, as shown in the first part of Tab. 1, with increases of +0.8 in BD, +0.6 in SBD and +0.19 in ||||DiC||||. Fig. 5 row 1 demonstrates that GMT is able to delineate the shape of small leaves at the plant centre more accurately.
MSU-PID. Tab. 1 second part shows that GMT surpass Mask2Former by +0.4 on BD and SBD, respectively, while performing on par in ||||DiC||||. Fig. 5 row 2 shows that GMT mitigates under-segmentation when one tiny leaf overlaps another leaf, whereas Mask2Former merges the two leaves and misses part of the outer shape; row 3 illustrates that GMT reduces the occurrence of missing leaves.
KOMATSUNA. As shown in the third section of Tab. 1, GMT improves over Mask2Former with a +0.5 increase in BD and slightly outperforms it in SBD by +0.1, while showing a minor underperformance in ||||DiC|||| by -0.02. Fig. 5 row 4 demonstrates that GMT avoids confusing similar green objects with leaves, and row 5, again, shows that GMT reduces under-segmentation in occlusion scenarios.
@@
The results shown in Tab. 3 demonstrate that for small and medium leaves on both datasets, GMT improve the segmentation quality (i.e. BD and SBD) notably while its counting ability remains on par with Mask2Former. For large leaves in MSU-PID, GMT performs worse in BD but outperforms in SBD and ||||DiC||||. As discussed in Sec. 4.2, BD may underestimate the errors caused by under-segmentation or missing leaves due to its one-directional matching nature, while SBD is a more balanced metric for segmentation quality. In fact, the better ||||DiC|||| score of GMT suggests that, despite a lower BD compared to Mask2Former, GMT is more effective at identifying all leaves. Finally, for large leaves in KOMATSUNA, GMT performs comparably to Mask2Former. Overall, GMT excels in segmenting small and medium leaves, which are more challenging in leaf instance segmentation, while maintaining solid performance on large leaves.